OpenStudy (toxicsugar22):

A classroom has a volume of 288m3 What is its volume in cm3?

sam (.sam.):

Well same thing, cube root 288 to get meters

OpenStudy (toxicsugar22):

0.006604

sam (.sam.):

Then convert meters to centimeters then cube the value

sam (.sam.):

6.604 meters to centimeters is?

OpenStudy (toxicsugar22):

would we have to multiple it threetimes

sam (.sam.):

Yes

OpenStudy (toxicsugar22):

660.4 centimeters

OpenStudy (toxicsugar22):

right

sam (.sam.):

Well done

sam (.sam.):

Now you want it cm3 so, cube it

OpenStudy (toxicsugar22):

288019036.9

OpenStudy (toxicsugar22):

right

sam (.sam.):

Um you should put it in scientific notation that'll make things cleaner, do you know how?

OpenStudy (toxicsugar22):

2.8801

OpenStudy (toxicsugar22):

right

sam (.sam.):

Yep, what are the exponent for 2.8801?

sam (.sam.):

\[\huge 2.8801 \times 10^{?}\]

OpenStudy (toxicsugar22):

8

sam (.sam.):

Good job

sam (.sam.):

We're done :)

OpenStudy (toxicsugar22):

can you help with one more please

sam (.sam.):

\[\Large 2.88\times 10^8 \text{cm}^3\]

sam (.sam.):

Okay I might have some time left

OpenStudy (toxicsugar22):

a NASA satellite showed that in 2009 the ozone hole over Antarctica had a maximum surface area of 24.1 million km2. The largest ozone hole on record occurred in 2006 and had a surface area of 29.6 million km 2. Calculate the difference in diameter (in meters) between the ozone hole in 2009 and in 2006.

sam (.sam.):

They want the 'difference', so what's your approach?

OpenStudy (toxicsugar22):

subtraction

sam (.sam.):

Mhm good

sam (.sam.):

Um we should not have negative areas, lol

OpenStudy (toxicsugar22):

yes so 29.6-24.1

OpenStudy (toxicsugar22):

right

sam (.sam.):

Yup

OpenStudy (toxicsugar22):

5.5

sam (.sam.):

Now, they say the values are in "million km2" which means 5.5 million km2, or \[5.5 \times 10^{6} km^2\]

OpenStudy (toxicsugar22):

the anwer is 5.5*10^6

sam (.sam.):

For km2 to km, you'll be square rooting \(5.5\times 10^{6}\)

sam (.sam.):

Nope, they want it in meters

sam (.sam.):

First you will convert km2 to km, then, multiply by 1000 for the conversion of km to m, then, square your answer to get m2

sam (.sam.):

Like this \[km^2 \rightarrow km \\ \sqrt{5.5 \times 10^6km^2}=2345km\] \[km \rightarrow m \\ 2345 km \times 1000=2345000m\] \[m \rightarrow m^2 \\(2345000m)^2=2345000^2m^2=5.5 \times 10^{12}m^2\]

sam (.sam.):

Yep, bob's ur uncle :)

OpenStudy (toxicsugar22):

can you do one more

sam (.sam.):

I gotta go, post a link in the math chat see if someone responds :) cya